OpenStudy (anonymous):

A swimming pool is 20 ft wide, 40 ft long, 3 ft deep at the shallow end, and 9 ft deep at is deepest point. If the pool is being filled at a rate of 1.4 ft3/min, how fast is the water level rising when the depth at the deepest point is 5 ft? (Round your answer to five decimal places.) Please Help! :P

OpenStudy (anonymous):

what math r u in?

OpenStudy (anonymous):

Calc 1

OpenStudy (anonymous):

have u done integration yet?

OpenStudy (anonymous):

Not yet, just finished learning about differentiation >< haha

OpenStudy (anonymous):

ok ill do this by differentiation then

OpenStudy (anonymous):

Its a related rates questions :(

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

v=volume h=height of water l=length the water makes w=width if u look at it sideways it would make a trapezoid and until the water reached 6 ft the volume would be v=1/2hlw

OpenStudy (anonymous):

now we need to v in terms of 1 variable so we need another equation

OpenStudy (anonymous):

Ok, do i differentiate V? or just plug in V=(1/2)[(40)(3+9)(20)]?

OpenStudy (anonymous):

now the width is a constant because we are only changing the vertical level of the water

OpenStudy (anonymous):

u have to differentiate v but we need to write the equation with only2 variables. v and h so the l is a problem

OpenStudy (anonymous):

what we are trying to find is the rate of change of the h when the height equals 5

OpenStudy (anonymous):

hmm ok :]

OpenStudy (anonymous):

ok the answer to that is 4,800

OpenStudy (anonymous):

now we need to find an equation that only uses l and h so that we can solve for l and write the volume equation in terms of h

OpenStudy (anonymous):

now the easiest way to do this is to look at similar traingles.

OpenStudy (anonymous):

if u drew it u would get a large triangle and a small triangle. the length of the large trianglee would b 40 or the length of the pool and the height of the large triangle would b 6(9-6) or how much deeper the pool is than the shallow end

OpenStudy (anonymous):

the small triangle would b length l and height h

OpenStudy (anonymous):

so we would have the formula: 40/6=l/h and we would solvefor h 40h/6=l so now we can substitue it back into our volume equation

OpenStudy (anonymous):

so v=(1/2)l*h*w becomes v=(1/2)(40/6)h^2*w

OpenStudy (anonymous):

(1/2)[(l/40)(12)(20)]

OpenStudy (anonymous):

oh nvm haha

OpenStudy (anonymous):

udont wanna put in the values yet because u want variables. only variables can be differentiated

OpenStudy (anonymous):

ok, so differentiate the (1/2)(40/6)h^2(w)?

OpenStudy (anonymous):

yes but simplify it first will make it easier

OpenStudy (anonymous):

5h^2*w

OpenStudy (anonymous):

also w is a constant so treat it as u would any number like 6 or something

OpenStudy (anonymous):

i got (10/3)h^2(w)

OpenStudy (anonymous):

oh ok haha

OpenStudy (anonymous):

my bad :)

OpenStudy (anonymous):

wait how did u get 5?

OpenStudy (anonymous):

i added 2 and 6 -_- instead of multiplying haha

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

dv/dt=(20/3)h?

OpenStudy (anonymous):

almost. u r missing 2 things

OpenStudy (anonymous):

dv/dt=dh/dt * dv/dh (20/3)h or something like that? ><

OpenStudy (anonymous):

well u got 1 of them. u r differentiating in terms of time or dt so that should b the onlydenominator. so the dv/dh is wrong but u r still missing 1 more thing

OpenStudy (anonymous):

umm...haha wait i have no idea :(

OpenStudy (anonymous):

well look at the equation for V that i gave u. what letter is in there thats not in ur derivative?

OpenStudy (anonymous):

w! but when you derive that wouldnt it become 1?

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

w is a constant. when u derive 6x in terms of it becomes 6 right? similarly if u derived wx in terms of x and w is a constan it would become w

OpenStudy (anonymous):

oh haha ok (20/3)hw then

OpenStudy (anonymous):

and then plug in the 5 so 100/3 (w)?

OpenStudy (anonymous):

(20/3)hw(dh/dt)

OpenStudy (anonymous):

oh then 1.4/ [(100/3)]w =dh/dt

OpenStudy (anonymous):

? hopefully ...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

isnt the w= 20?

OpenStudy (anonymous):

u r dividing the walso right?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

I got 0.0021?

OpenStudy (anonymous):

(1.4)(3/100)(1/w) will give u ur answer

OpenStudy (anonymous):

ok thank you!! :D